OpenStudy (kaliyah0524):

Solve the proportion for x 1+3x/4 = 5/2

OpenStudy (dayakar):

r u there

OpenStudy (kaliyah0524):

Yes

OpenStudy (dayakar):

subtract both sides 1

OpenStudy (michele_laino):

I agree with @dayakar , so I can write this: \[1 + \frac{{3x}}{4} - 1 = \frac{5}{2} - 1\] please simplify

OpenStudy (kaliyah0524):

1+2/4=4/2 . ?

OpenStudy (dayakar):

3x/4 = 2.5 -1 ( 5/2 =2,5) 3x/4 = 1.5

OpenStudy (michele_laino):

hint: left side: \[1 + \frac{{3x}}{4} - 1 = 1 - 1 + \frac{{3x}}{4} = 0 + \frac{{3x}}{4} = ...?\]

OpenStudy (michele_laino):

right side: \[\frac{5}{2} - 1 = \frac{{5 - \left( {1 \times 2} \right)}}{2} = \frac{{5 - 2}}{2} = ...?\]

OpenStudy (dayakar):

\[\frac{ 4 }{ 3}*\frac{ 3x }{ 4 }= 1.5 *\frac{ 4 }{ 3}\]

OpenStudy (welshfella):

could the question be